### v1

# Rigid-Flexible Values for Embeddings of Four-Dimensional Ellipsoids into Almost-Cubes

Abstract.
We consider the embedding function $c_{b}(a)$ describing the problem of symplectically embedding an ellipsoid $E(1,a)$ into the smallest possible scaling by $\lambda>1$ of the polydisc $P(1,b)$. In particular, we calculate rigid-flexible values, i.e. the minimum $a$ such that for $E(1,a^{\prime})$ with $a^{\prime}>a$, the embedding problem is determined only by volume. For $21<b<$ 2 we find that these values vary piecewise smoothly outside a discrete set of discontinuities at $b\in\left(\frac{n+1}{n}\right)^{2}$.
Key words and phrases:
Differential geometry, symplectic geometry MSC codes: 53D05, 53D22
1. Introduction, Statement of Results
The problem of embedding one symplectic manifold into another touches on a wide variety of topics in symplectic geometry, and in this work we focus in particular on embeddings of ellipsoids $E(a,b)$ into polydiscs $P(a,b)$ Here, a polydisc
$P(a,b):=\left\{(z_{1},z_{2})\in\mathbb{C}^{2}\,|\,\pi\|z_{1}\|^{2}<a,\pi\|z_{2}\|^{2}<b\right\}$
is the 4-dimensional open symplectic manifold $B_{2}(a)\times B_{2}(b)\subset\mathbb{C}^{2}$, where each factor is a 2-disc of fixed radius centered at $0\in\mathbb{C}$. Similarly the ellipsoid $E(a,b)$ is given by
$E(a,b):=\left\{(z_{1},z_{2})\in\mathbb{C}^{2}\,|\,\frac{\pi\left|\left|z_{1}\right|\right|^{2}}{a}+\frac{\pi\left|\left|z_{2}\right|\right|^{2}}{b}<1\right\}.$
We study the problem of finding embeddings of an ellipsoid into a polydisc. This information is encoded by a function $c_{b}(a)$ whose value at $a$ is the smallest $\lambda$ such that
(1.1) $E(1,a)\hookrightarrow P(\lambda,\lambda b).$
More precisely,
$c_{b}(a):=\inf\{\lambda>0\,|\,E(1,a)\hookrightarrow P(\lambda,\lambda b)\}.$
Earlier work on embeddings of ellipsoids into polydiscs [2, 3] considered a fixed value of $b$ and calculated $c_{b}(a)$ in its entirety for all $a$. For example, $c_{1}(a)$ contains a so-called “infinite staircase,” a convergent sequence of $a$-values $a_{n}\to a_{\infty}$ such that $c_{b}(a)$ is non-decreasing and linear or constant on each $[a_{i},a_{i+1}]$. In addition, they find that for sufficiently large values of $a$, $c_{b}(a)$ coincides with the volume constraint $\sqrt{\frac{a}{2b}}$: for sufficiently elongated ellipsoids the only obstruction to the existence of symplectic embeddings is the volume constraint. In later work [2, 10] both the source and target in the embedding problem were allowed to become elongated, with the goal of analyzing the resulting changes in the embedding function $c_{b}(a)$.
To set notation, define
$RF(b):=\inf\left\{A\,\,|\,\,c_{b}(a)=\sqrt{\frac{a}{2b}}\,\,\mbox{for all}\,\,a\geq A\right\}.$
We call this number the rigid-flexible value, or the RF-value.
Varying this extra parameter $b$ uncovers delicate structure in the the $b$-direction for functions $c_{b}(a)$. In [2] the authors find that for $b\in\mathbb{N}$ with $b\geq 2$, $c_{b}(a)$ contains no infinite staircases and can be computed explicitly for all $a$. On the other hand, [10] provides examples of infinite sequences (some arbitrarily large and others converging to 1) of irrational $b$ such that for each such $b$, $c_{b}(a)$ contains an infinite staircase. [10] also computes the embedding functions $c_{b}(a)$ for $b\in[1,\infty)$ and $a\in[1,3+\sqrt{2}]$. Previous joint work by the second author computed $RF(b)$ in this problem for $b>2$, finding that it was piecewise smooth in $b$ away from an infinite discrete set of discontinuities.
In this paper we examine the same problem but for $21<b<$ 2. We term these polydiscs “almost-cubes” following [3], where cubes are polydiscs $P(\lambda,b\lambda)$ with $b=1$. In this region we again see an infinite set of discontinuities, this time accumulating at 1, but away from these points $RF(b)$ is smooth.
Theorem 1.1.
Let $b\in(1,2)\setminus\left\{\frac{(n+1)^{2}}{n^{2}}\right\}$ for $n\geq 3$. Then for $k=\lfloor\frac{n}{2}\rfloor$,
(1.2) $RF(b)=\begin{cases}2b\left(\frac{8}{2k^{2}+k+b(2k^{2}-k)}\right)^{2}&\mbox{when $n$ is even}\\ 2b\left(\frac{8k^{2}+8k+1}{(2k+1)(k+1)+b(2k^{2}+k)}\right)^{2}&\mbox{when $n$ is odd}\\ \end{cases}$
determined by the obstructive classes (for $k\geq 1$)
$S_{k}$ $:=((2k+1)(k+1),k(2k+1);(k^{2}+k+1),(k^{2}+k)^{\times 7}),$
$T_{k}$ $:=(2k^{2}+k,2k^{2}-k;(k^{2})^{\times 7},k^{2}-1).$
As an example, for $b\in\left(\frac{16}{9},2\right)$ we have $n=3$ so we consider the $S_{k}$-classes with $k=1$. In this interval, the class $S_{1}=(6,3;3,2^{\times 7})$ is obstructive at $a=8$. Then, for $b\in\left(\frac{25}{16},\frac{16}{9}\right)$, we have $n=4$ so we consider the $T_{k}$-classes with $k=2$. In this interval the class $T_{2}=(10,6;4^{\times 7},3)$ is obstructive at $a=8$. Then for $b\in\left(\frac{36}{25},\frac{25}{16}\right)$ we find $S_{2}=(15,10;7,6^{\times 7})$ is obstructive at the same $a=8$.
Consequently, the behavior of $RF(b)$ on $21<b<$ 2 differs from the left continuity which appears for $b>2$, calculated in [12]. In this case $RF(b)$ is neither left nor right continuous. Experimental evidence suggests that $RF(b)$ at these discontinuities is determined by obstructive classes at $a=7$. This resembles the case $b=1$, for which [3] shows that $RF(1)=7\frac{1}{32}$ whereas [2] shows that $RF(2)=8\frac{1}{36}$.
1.1. Outline of Paper
In Section 3 we first describe the exceptional classes which determine the $RF$-value for $b\not\in\left\{\left(\frac{n+1}{n}\right)^{2}\right\}$. Then, in Section 4, we describe the computer-assisted search establishing that no more obstructive classes can appear when $a\in(8,9]$. The remainder of the paper is devoted to applying the reduction method for $a$ in the interval $[9,\infty)$, to establish that there are no obstructions to embeddings other than the volume constraint. This argument splits into many cases, according to the ordering of the terms.
1.2. Acknowledgments
The second author thanks Dan Cristofaro-Gardiner for suggesting this problem and providing support and encouragement along the way.
This work used the NCSA Delta CPU at the University of Illinois Urbana-Champaign through allocation MTH230018 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
2. Preliminaries: Two Methods for Finding Symplectic Embeddings
Here, we review the methods we use for detecting symplectic embeddings. This is not an exhaustive list, and more detailed expositions are in [2, 9], so we review only what we use. The following definition is central to both methods. Fix $b\geq 1$. Since the function $c_{b}(a)$ is continuous in $a$, it suffices to compute it for $a\geq 1$ rational.
Definition 2.1.
The weight expansion $\textbf{w}(a,b)$ of such an $a$ is the finite decreasing sequence
$\textbf{w}(a)=(1^{\times\ell_{0}},w_{1}^{\times\ell_{1}},...,w_{n}^{\times\ell_{n}}),$
where $w_{1}=a-\ell_{0}<1,w_{2}=1-\ell_{1}w_{1}<w_{1}$, and so on.
We write $\textbf{w}(a)$ to mean $\textbf{w}(1,a)$, and we write $w(a_{i})$ to mean the $i$th term of the weight expansion $\textbf{w}(a)$.
In [7, Thm. 1.1], it is shown that there is a canonical decomposition of any ellipsoid $E(1,a)$ into a collection of balls
$E(1,a):=\coprod_{i}B(w_{i})$
where the $w_{i}$ are the terms in the weight expansion $\textbf{w}(a)$.
Theorem 2.2 ([3]).
Let $a,b\in\mathbb{Q}$ be positive. There exists a symplectic embedding $E(1,a)\hookrightarrow P(\lambda,\lambda b)$ if and only if there is a symplectic embedding
$E(1,a)\coprod B(\lambda)\coprod B(\lambda b)\hookrightarrow B(\lambda(b+1)).$
This reduces the polydisc problem to a ball-packing problem of embedding balls of capacity $e_{i}$ into a ball of capacity $\mu$:
(2.1) $\coprod_{i}B(e_{i})\hookrightarrow B(\mu).$
The purpose of introducing these constructions is that a collection of balls can be embedded symplectically precisely when the associated multiple blow-up of $\mathbb{C}P^{2}$ carries a symplectic form in a certain cohomology class. We denote by $\overline{\mathcal{C}}_{K}(X_{n})$ the set of cohomology classes represented by symplectic forms for which the anticanonical class is $K=-3L+\sum_{i}E_{i}$, where $L$ is Poincaré dual to a line in $\mathbb{C}P^{2}$ and each $E_{i}$ is dual to the $i$th exceptional sphere. By [8], the embedding (2.1) exists when the following cohomology class is in the symplectic cone:
$\mu L-\sum_{i}e_{i}E_{i}\in\overline{\mathcal{C}}_{K}(X_{n}).\newline \indent$
The above fact gives a sufficient criterion for a class to lie in $\overline{\mathcal{C}}_{K}(X_{n})$. If there is a symplectic form in a given class, it must have non-negative intersection with certain holomorphic curves. By [6], this is also sufficient: if $\mathcal{E}_{K}(X_{n}):=\{e\in H_{2}(X_{n})\,|\,\langle e,e\rangle=-1,\,\,\langle K,e\rangle=-1\}$ is the set of exceptional classes, then we may characterize the symplectic cone referenced above as
(2.2) $\overline{\mathcal{C}}_{K}(X_{n})=\{\omega\in H^{2}(X_{n})\,|\,\langle\omega,e\rangle\geq 0\,\,\forall e\in\mathcal{E}_{k}(X_{n})\}.$
In summary, an embedding of the form (2.1) exists when the associated weight expansion represents a cohomology class in the symplectic cone $\overline{\mathcal{C}}_{K}(X_{n})$, and the equality (2.2) characterizes elements of the symplectic cone as those which pair non-negatively with so-called exceptional classes $\mathcal{E}_{k}(X_{n})$.
We now require a condition for determining when homology classes are exceptional. With respect to the basis of $H^{2}(X_{n};\mathbb{R})\simeq\mathbb{R}^{n+1}$ given above, a Cremona transform is the map given by
$(d;m_{1},...m_{n})\mapsto(2d-m_{1}-m_{2}-m_{3};d-m_{2}-m_{3},d-m_{1}-m_{3},d-m_{1}-m_{2},m_{4},...,m_{n})$
We will use Cremona transforms in two different contexts below. First, we can state the condition for an integral homology class to lie in $\mathcal{E}_{k}$, which is proven in [9] based on work of [5, 6].
Theorem 2.3.
A class $(d;m_{1},...,m_{n})\in H_{2}(X_{n};\mathbb{Z})$ is in $\mathcal{E}_{K}(X_{n})$ if and only if its entries satisfy the Diophantine equations
[…]
$d^{2}+1=\sum_{i}m_{i}^{2}$
and $(d;m_{1},...,m_{n})$ reduces to $(0;-1,0,\ldots,0)$ after a sequence of Cremona transforms.
For the problem of embedding ellipsoids into polydiscs, the natural compactification of $P(a,b)$ adds a single point to each disc, yielding $S^{2}\times S^{2}$. This manifold is in fact diffeomorphic to the 2-fold blow-up of $\mathbb{C}P^{2}$, so the $n$-fold blow-up of $S^{2}\times S^{2}$ (denoted $Y_{n}$) can be identified with $X_{n+1}$. The induced isomorphism on cohomology $\psi:H^{2}(Y_{n})\to H^{2}(X_{n+1})$ is given by
$(d,e;m_{1},m_{2},m_{3},...,m_{n})\mapsto(d+e-m_{1};d-m_{1},e-m_{1},m_{2},...,m_{n})$
Using this isomorphism we may translate the Diophantine equations (2.3) into the following conditions.
[…]
(2.5) $\sum_{i}m_{i}^{2}$ $=$ $1\displaystyle 2de+$ 1
2.1. Obstructive Classes
We can now describe the first of the two methods we employ to find symplectic embeddings. If $A:=(d,e;\textbf{m})\in\mathcal{E}_{k}$ with the basis given previously, the condition
(2.6) $\lambda=\frac{\langle\textbf{w}(a),\textbf{m}\rangle}{d+be}>0$
is equivalent to
(2.7) $\lambda(d+be)-\langle\textbf{w}(a),\textbf{m}\rangle>0.$
The following statement is [2, Method 1’].
Theorem 2.4.
An embedding (1.1) exists iff $\lambda\geq\sqrt{\frac{a}{2b}}$ and
$\mu_{b}(d,e;m):=\frac{\sum_{i}m_{i}\cdot w_{i}(a)}{d+be}\leq\lambda$
for every $(d,e;m_{1},...,m_{n})\in H_{2}(Y_{n};\mathbb{Z})$ which satisfies equations (2.4, 2.5) and reduces to $(0;-1)$ after some sequence of Cremona transforms.
Definition 2.5.
We say that a class $A$ is obstructive at $a>0$ if the quotient in (2.6) is larger than the volume constraint $\sqrt{\frac{a}{2b}}$.
To use this method, then, we must find all obstructive classes. In practice, this method proceeds first by finding a discrete subset of $a$ which have a special relationship to the obstructive classes, described as follows.
If the graph of $c_{b}(a)$ does not follow the volume constraint, then locally it must be given by the obstruction function of some class $(d,e;\textbf{m})$. Restricting to the interval where this class determines the graph, [9] Lemma 2.1.3 states that there is a particular $a$-value on the interval whose weight expansion coincides with the number of positive entries of m, the tail of the obstructive class.
Lemma 2.6.
Let $C=(d,e;\textbf{m})$ be an exceptional class, and let $I$ be the maximal nonempty open set on which $\mu_{a,b}(C)>\sqrt{\frac{a}{2b}}$ for all $a\in I$. Then, there is a unique element $a_{0}\in I$ such that the length of the weight expansion for $a_{0}$, denoted $\ell(a_{0})$, satisfies $\ell(a_{0})=\ell(\textbf{m})$, and moreover $\ell(a)>\ell(\textbf{m})$ if $a\in I\setminus\{a_{0}\}$.
The following quantities will be relevant going forward: let $\ell_{0}$ denote the number of $1$’s in the weight expansion of $a$ and subsequently let $\ell_{i}$ denote the lengths of subsequent blocks. When a class $(d,e;\textbf{w}(a))$ is obstructive, we have a vector of error terms $\epsilon$ defined as
$\textbf{m}=\frac{d+be}{\sqrt{2a}}\textbf{w}(a)+\epsilon.$
Each contribution to this error, thought of as the difference between m and $\frac{d+be}{\sqrt{2a}}\textbf{w}(a)$, will be important, so we define $v_{i}=\frac{d+be}{\sqrt{2a}}w_{i}$ for $i=0,...,M$ where $M$ denotes the length of the weight expansion of $a$.
Also let
$\sigma=\sum_{\ell_{0}+1}^{M}\epsilon_{i}$
With these terms established, the following result (proven in [2, 9]) will be used repeatedly.
Lemma 2.7.
For all $(d,e;\textbf{m})\in\mathcal{E}$, suppose that $a\in\mathbb{Q}$ and $b\in\mathbb{R}$ with $a,b>1$.Then we have
  1. (1)
    If $\mu_{b}(d,e;\textbf{m})(a)>\sqrt{\frac{a}{2b}}$ then $d=be+h$ where $|h|<\sqrt{2b}$ and $\langle\epsilon,\epsilon\rangle<1-\frac{h^{2}}{2b}$
The second step of the obstructive class method establishes that in fact only finitely many classes must be checked at the values of $a$ characterized in Lemma 2.6. This is the purpose of the inequality (2.9) below. Proofs of this fact appear in [9, 2]. A proof of the inequality (2.9) appears in [12].
Lemma 2.8.
Let $(d,e;\textbf{m})$ be an exceptional class such that $\ell(a)=\ell(m)$ and $\mu_{b}(d,e;\textbf{m})>\sqrt{\frac{a}{2b}}$, and $b\in[1,2]$. Set $v_{M}:=\frac{(d+be)\sqrt{2b}}{q(b+1)\sqrt{a}}$ where $q$ is the last denominator in the weight expansion $\textbf{w}(a)$. Then
  1. (1)
    If $v_{M}\leq\frac{1}{2}$, then $v_{M}>\frac{1}{3}$ and $\sigma^{\prime}\leq\frac{1}{2}$. If $v_{M}\leq\frac{2}{3}$, then $\sigma^{\prime}\leq\frac{7}{9}$. with $\delta=y(a)-\frac{1}{q}$ and $y(a)=a+1-2\frac{b+1}{\sqrt{2b}}\sqrt{a}$, we have
    (2.8) $\sqrt{q+\left\lfloor a\right\rfloor+1}\geq 1+\delta v_{M}q$
    One can also show the bound from [2, 3], which is
    (2.9) $2be+h\leq\frac{\sqrt{2ba}}{\delta}\left(\sqrt{\sigma q}-(1-h(1-\frac{1}{b})\right)\leq\frac{\sqrt{2ba}}{\delta}\left(\frac{\sigma}{\delta v_{M}}-\left(1-h(1-\frac{1}{b})\right)\right)$
The following lemma, [9, Lem. 2.1.8] will also be necessary for our arguments.
Lemma 2.9.
Assume that $(d,e;\textbf{m})$ is an exceptional class such that $\mu(d,e;\textbf{m})>\sqrt{\frac{a}{2b}}$ for some $a$. Let $J:=\{k,...,k+s-1\}$ be a block of $s-1$ consecutive integers for which $w(a_{i}),i\in J$ is constant. Then $(m_{1},...,m_{k+1})$ is of the form
[…]
$(m,...,m+1).$
Moreover, there is at most one block of length $s\geq 2$ on which the $m_{i}$ are not all equal, and if $m_{1}\neq m_{k+1}$, then $\sum_{i=k}^{k+s-1}\epsilon_{i}^{2}\geq\frac{s-1}{s}$.
2.2. Reduction at a Point
Though the first method is a necessary and sufficient condition for the embedding of an ellipsoid into a polydisc, in principle one might need to check more exceptional classes than is computationally feasible. The following method straightforwardly determines whether an embedding exists but the complexity of the process depends strongly on the value of $a$.
Definition 2.10.
The defect $\delta$ of an ordered vector $(d;m_{1},...,m_{n})$ is the sum $d-m_{1}-m_{2}-m_{3}$.
The following is established in [1, 4].
Theorem 2.11.
An embedding $E(1,a)\hookrightarrow P(\lambda,\lambda b)$ exists if there exists a finite sequence of Cremona moves that transforms the ordered vector
$((b+1)\lambda;b\lambda,\lambda,\textbf{w}(a))$
to an ordered vector with non-negative entries and defect $\delta\geq 0$.
We will apply this repeatedly in the proofs of both Theorem 1.1 and Theorem 1.2.
Notice the requirement that the Cremona moves be performed on ordered vectors. The following fact [2, Prop. 2.2] allows us to avoid re-ordering.
Proposition 2.12.
Let $\alpha=(\mu;a_{1},...,a_{n})$ be a vector with $\mu\geq 0$ and $\alpha^{2}\geq 0$ and assume that there is a sequence $\alpha=\alpha_{0},...,\alpha_{m}$ of vectors such that $\alpha_{j+1}$ is obtained from $\alpha_{j}$ by a sequence of Cremona moves. If $\alpha_{m}=(\hat{\mu};\hat{\alpha}_{1},...,\hat{\alpha}_{n})$ is reduced and $\hat{\alpha}_{1},...,\hat{\alpha}_{n}\geq 0$, then $\alpha\in\overline{\mathcal{C}}_{K}(X_{n})$.
Thus to find embeddings we need only prove enough about the ordering to ensure that the defect at a certain step is non-negative, provided that we have non-negativity of the terms in each vector. So, only knowledge of the ordering for the first three terms is strictly necessary. We will usually verify the ordering at each step, except when doing so results in distinguishing too many cases.
3. Classification of Exceptional Classes at $a=8.$
In this section, we narrow the search for exceptional classes at $a=8$ and for all $1<b\leq 2$ by considering the constraints that Equations 3.3 and 3.4 put on candidate vectors of the form $\langle d,e;\mathbf{m}\rangle.$ For each $n\geq 0,$ recall that we have
(3.1) $S_{k}:=\langle 2k^{2}+k,2k^{2}-k;(k^{2})^{\times 7},(k^{2}-1)\rangle,$
and let
(3.2) $T_{k}:=\langle(2k+1)(k+1),k(2k+1);(k^{2}+k+1),(k^{2}+k)^{\times 7}\rangle.$
We seek to prove the following result:
Lemma 3.1.
For $1<b\leq 2,$ the only exceptional classes $\langle d,e;\mathbf{m}\rangle$ with $\mu_{b}(d,e;\mathbf{m})(8)>\sqrt{8/2b}$ are $S_{k}$ and $T_{k}$ for some $n\geq 0.$
To begin, recall that if $\langle d,e;\mathbf{m}\rangle$ is obstructive, where $\mathbf{m}=(m_{1},\ldots,m_{k}),$ then
(3.3) $2de+1=\sum_{i=1}^{k}m_{i}^{2},$
and
(3.4) $2(d+e)-1=\sum_{i=1}^{k}m_{i}.$
We also make use of the fact that if $a$ is an integer, then the shape of $\mathbf{m}$ is well-understood:
Lemma 3.2.
If $\langle d,e;\mathbf{m}\rangle$ is obstructive at $a=8,$ then $\mathbf{m}=(m^{\times 8}),(m^{\times 7},m-1)$ or $(m+1,m^{\times 7}).$
The next number-theoretic lemma will be needed to calculate values of $d,e$ explicitly.
Lemma 3.3.
Let $x\in\mathbb{Z}.$ Then $8|(x^{2}-1)$ if and only if $x=\pm 1\pmod{4}.$
3.1. Proof of Lemma 3.1
By [10, Eqn 4.2, p.48], we may assume $d\geq e.$ Suppose $\langle d,e;\mathbf{m}\rangle$ is obstructive at $a=8.$ If $\mathbf{m}=(m^{\times 8}),$ then $2de+1=8m^{2}$ by Equation 3.3, which is a contradiction because $d,e$ and $m$ are all integers. Therefore, the only two cases we need to consider are $\mathbf{m}=(m^{\times 7},m-1)$ and $\mathbf{m}=(m+1,m^{\times 7}).$
3.2. Case I: $\mathbf{m}=(m^{\times 7},m-1)$
In this case, we have:
(3.5) $2de+1=8m^{2}-2m+1,$
and
(3.6) $12(d+e)-1=8m-$ 1 .
By Equation 3.6, $m=(d+e)/4.$ Substituting $m$ into Equation 3.5, we see that
(3.7) $de=\frac{(d+e)^{2}}{4}-\frac{d+e}{4}.$
Let $h=d+e.$ Then Equation 3.7 becomes
(3.8) $2h=1+4e\pm\sqrt{8e+1}.$
Since $h\in\mathbb{Z},$ it follows that $18e+$ 1 is a perfect square. In particular, if $x^{2}=8e+1$ for some $x\in\mathbb{Z},$ then, by Lemma 3.3, $x=\pm 1+4n$ for some $n\in\mathbb{Z}.$ Thus, either
(3.9) $e=2n^{2}+n,$
or
(3.10) $e=2n^{2}-n.$
[…]
$d=h-e=2n^{2}+n.$
If $e=2n^{2}+n,$ then $8e+1=(4n+1)^{2}$, and like before we have $h=4n^{2}$ so that $d=2n^{2}-n.$ However, since we are assuming that $d\geq e,$ this solution is excluded.
In summary, if $\langle d,e;\mathbf{m}\rangle$ is obstructive at $a=8,$ where $\mathbf{m}=\langle m^{\times 7},m-1\rangle,$ then
$\langle d,e;\mathbf{m}\rangle=\langle 2n^{2}+n,2n^{2}-n;(n^{2})^{\times 7},(n^{2}-1)\rangle.$
Since $d\geq e,$ we have $2n^{2}+n\geq 2n^{2}-n$ which implies $n\geq 0.$
3.3. Case II: $\mathbf{m}=(m+1,m^{\times 7})$
By Equations 3.3 and 3.4, we have:
(3.11) $de=4m^{2}+m,$
and
(3.12) $d+e=4m+1.$
By Equation 3.11, $de=m(4m+1)=m(d+e).$ Solving for $m$ in Equation 3.12 and substituting the result into the latter equation, we see that
(3.13) $4de=(d+e-1)(d+e).$
Let $h=d+e.$ Rewrite Equation 3.13 in terms of $h$ to obtain
(3.14) $h^{2}-(1+4e)h+4e^{2}=0.$
Therefore,
(3.15) $2h=1+4e\pm\sqrt{1+8e},$
which is exactly Equation 3.8.
Therefore, we have $e=2n^{2}+n,$ or $e=2n^{2}-n.$ We conduct our analysis in a way completely analogous to what we did in Case I. Suppose $e=2n^{2}+n.$ Then $h=4n^{2}$ or $h=4n^{2}+4n+1.$ In this scenario, the first case is impossible because $d+e=1\pmod{4}.$ Therefore,
$d+e=4n^{2}+4n+1,$
and thus
(3.16) $d=4n^{2}+4n+1-e=2n^{2}+3n+1=(2n+1)(n+1).$
Just like above, the case where $e=2n^{2}-n$ is excluded because $d\geq e.$ Now, $m=n^{2}+n$ by Equation 3.12, so we see that if $\langle d,e;\mathbf{m}\rangle$ is obstructive at $a=8$ for some $b\geq 1,$ where $\mathbf{m}=(m+1,m^{\times 7}),$ then we must have
$\langle d,e;\mathbf{m}\rangle=\langle(2n+1)(n+1),n(2n+1);(n^{2}+n+1),(n^{2}+n)^{\times 7}\rangle$
for some $n\in\mathbb{Z}.$ As before, since $d\geq e,$ we have $2n^{2}+3n+1\geq 2n^{2}+n$ which implies $n\geq 0.$
4. No obstructive classes for $a\in(8,9]$
In this section we outline the computer-assisted search method used to investigate the interval $(8,9)$. Using the equation (1) $d=be+h$ from of the same lemma, we find that for each $e$ the range of possible $d$ which might appear in a candidate obstructive class ranges from $22e-$ 2 to $22e+$ 2. Using the inequality (2.9) from Lemma 2.7, we find that the largest $q$ to check is 11, corresponding to $e\leq 40$, giving rise to 934 possible $(d,e)$-pairs.
It then remains to verify that no classes with these values of $(d,e)$ are obstructive. To accomplish this task, we perform a search along the lines of what appears in [9, Appendix B]. Our search is significantly less economical (to an extreme degree) but it does not depend at all of the value of $b$.
For fixed $(d,e)$ the self-intersection and Chern number conditions in equations (2.4) and (2.5) constrain the possible m which can appear. For example, by (2.5), the first term $m_{1}$ in m can be no larger than $\lfloor\sqrt{2de+1}\rfloor$. If $m_{1}$ takes this maximal value, the subsequent term can be no larger than
$\left\lfloor\sqrt{2de+1-\lfloor\sqrt{2de+1}\rfloor^{2}}\right\rfloor.$
This recursive structure suggests a straightforward method for enumerating candidate obstructive classes. Pseudocode for an algorithm Extend() carrying this out appears in (1), along with an example recursion tree. Computer code (written in Python) used to perform this search is available on GitHub. While this method will clearly produce all possible classes eventually, it has the significant downside of a high time complexity. Indeed, we may estimate an upper bound for this runtime in terms of $e$. Using $d\leq 2e+2$ and the right-hand sides of (2.4) and (2.5), the upper bound for the $\ell^{2}$ norm as an input of Extend() is $12e^{2}+$ 1. As the loop in line 15 of (1) may call the function $\lfloor\sqrt{2e^{2}+1}\rfloor$ times, this worst-case time complexity is $\mathcal{O}(e^{2e^{2}+1})$.
Fortunately, there are improvements possible on the version described above and shown in Algorithm (1). Lemma 2.6 aids in computation by obviating the need to calculate past the bound on $q$ from Lemma 2.7. (1) of this lemma bounds $q$ by 11, and the weight expansion of any number of the form $8\frac{n}{11}$ has length at most 19, so there is no need to to continue a search past this point. Even so, given the bound $e\leq 40$, the time and memory demands of the method necessitate high-performance computing resources both to generate possible classes and check them for possible obstructions. The additional memory available allows for the use of memoization, caching previously computed values rather than re-computing them.
After generating all possibilities, we use Lemma 2.9 to eliminate those with more than one block on which the entries differ by more than 1. As this depends on the weight sequence of $a$, we verify this for each of the possible $a$-values.
For an example computation, consider the class $(42,22;16,15^{\times 7},3,3)$. While this satisfies the conditions of Lemma 2.9 and Lemma 2.6 at $a=8\frac{1}{2}$, we can easily calculate that for any $b\in(1,2)$,
$\frac{\langle(16,15^{\times 7},3,3),(1^{\times 8},\frac{1}{2},\frac{1}{2})\rangle}{42+b\cdot 22}<\frac{\langle(16,15^{\times 7},3,3),(1^{\times 8},\frac{1}{2},\frac{1}{2})\rangle}{42+2\cdot 22}<\sqrt{\frac{8\frac{1}{2}}{2\cdot 1}}<\sqrt{\frac{8\frac{1}{2}}{2\cdot b}}$
Data: $\ell^{2}$ = remaining $\ell^{2}$-norm, $\ell^{1}$ = remaining $\ell^{1}$-norm, $p$ = previous term
Result: List of tuples Extensions=$\textbf{m}_{k}$}
$\textbf{Extensions}\leftarrow\{\}$;
if $\ell^{2}<\ell^{1}$ OR $\ell^{2}<0$ OR $\ell^{1}<0$ then
       return Extensions
end if
if $\ell^{2}$=$\ell^{1}$ then
       return $\{(1^{\times\ell^{2}})\}$
end if
if $p\neq 0$ then
       start = min($p$, $\sqrt{\ell^{2}})$
end if
else
       start $=\sqrt{\ell^{2}}$ ;
end if
for $i\leftarrow$ start to $1$ by -1 do
[…]
       end foreach
end for
[…]
Algorithm 1 Extend: a recursive algorithm for solutions to Equations (2.4) and (2.5)
5. Reduction Calculations
We follow the notational convention in [12]. In this sectoin we use Cremona reductions to show that the volume constraint is sufficient to guarantee embeddings if $a\geq 9$ and $1\leq b\leq 2$ Throughout, we write
$C_{k}=(h||t_{1},t_{2},t_{3},\ldots)$
if $C_{k}$ is an unordered weight vector, where the ordering of $t_{1},t_{2},\ldots$ is not necessarily known in full. If $C_{k}$ is an ordered weight vector, we write
$C_{k}=(h;t_{1},t_{2},t_{3},\ldots),$
and we let
(5.1) $\delta_{k}=h-t_{1}-t_{2}-t_{3}$
be the defect of $C_{k}.$
Let $a\geq 9,$ and let
$\mathcal{W}(a)=(1^{\times\left\lfloor a\right\rfloor},d_{1}^{\times l_{1}},\ldots,d_{k}^{\times l_{k}})$
be the weight expansion of $a.$ Let $1\leq b\leq 2$
Note that
(5.2) $d_{1}=a-\left\lfloor a\right\rfloor.$
Recall that the initial unordered vector is of the form
(5.3) $C_{0}=((b+1)\lambda||b\lambda,\lambda,\mathcal{W}(a))=((b+1)\lambda||b\lambda,\lambda,1^{\times\left\lfloor a\right\rfloor},d_{1}^{\times l_{1}},\ldots,d_{t}^{\times l_{t}}).$
Now $b\lambda\geq\lambda$ because $b\geq 1.$ Since $a\geq 9,$ we have $b\lambda>2,$ so it follows that
(5.4) $b\lambda-1>1.$
Either $\lambda\geq 2$ or not. If $\lambda\geq 2,$ then of course
$C_{0}=((b+1)\lambda;b\lambda,\lambda,1^{\times\left\lfloor a\right\rfloor},\ldots),$
with defect $\delta_{0}=-1,$ and so after performing a Cremona transformation, we obtain
$C_{1}=((b+1)\lambda-1;b\lambda-1,\lambda-1,1^{\times\left\lfloor a\right\rfloor-1},\ldots),$
with defect $\delta_{1}=0.$ Therefore, we assume
(5.5) $b\lambda-1>1>\lambda-1,$
and
$C_{1}=((b+1)\lambda-1||b\lambda-1,1^{\times\left\lfloor a\right\rfloor-1},\lambda-1,d_{1}^{\times l_{1}},\ldots),$
where the only remaining ambiguity in the ordering of the first several terms of $C_{1}$ is whether $\lambda-1\geq d_{1}$ or $\lambda-1<d_{1}.$ Both cases can occur for suitable $a$ and $b,$ and we address these cases below.
5.1. Case I: $\lambda-1\geq d_{1}$
In this case, we have
(5.6) $b\lambda-1>1>\lambda-1\geq d_{1}$
and thus
$C_{1}=((b+1)\lambda-1;b\lambda-1,1^{\times\left\lfloor a\right\rfloor-1},\lambda-1,d_{1}^{\times l_{1}},\ldots),$
with defect $\delta_{1}=\lambda-2.$
For each $k\in\mathbb{Z},$ define
(5.7) $h(k)=b\lambda-1+k(\lambda-2).$
Lemma 5.1.
For each $k\in\mathbb{Z},$ we have:
  1. (1)
    $h(k+1)=h(k)+\lambda-2;$
  2. (2)
    $h(k)\geq 1\implies h(k+1)\geq\lambda-1;$ and
  3. (3)
    $h(k)+3-2\lambda=h(k-2)-1.$
Proof.
(1) is clear; (2) follows by adding $\lambda-2$ to both sides; and for (3) we have
$h(k)+3-2\lambda$ $=$ $h(k)+(4-2\lambda)-1$
$=$ $(h(k)+(2-\lambda))+(2-\lambda)-1$
$=$ $(h(k-1)+2-\lambda)-1$
$=$ $h(k-2)-1.$
As an immediate and important consequence of Lemma 5.1, we have
$h(0)>h(1)>h(2)>\ldots$
Lemma 5.2.
Let $k\geq 0$ be an integer. If $h(k)\geq 1$ and $\left\lfloor a\right\rfloor-(2k+1)\geq 0,$ then
$C_{k+1}=(h(k)+\lambda;h(k),1^{\times\left\lfloor a\right\rfloor-(2k+1)},(\lambda-1)^{\times 2k+1},d_{1}^{\times l_{1}},\ldots).$
Proof.
$C_{1}$ has the stated form by 5.6 with defect $\delta_{1}=\lambda-2.$ The claim then follows by induction on $k$ combined with the identities of Lemma 5.1.∎
Throughout this section, it will be useful to know when, for positive real numbers $p,q,$ we have $(b+p)\lambda-q\geq 0$ for all $b\in[1,2].$ Now, $(b+p)\lambda-q\geq 0\iff a\geq 2q^{2}b/(b+p)^{2}.$ A straightforward analysis of the function $f(x)=2q^{2}x/(x+p)^{2}$ on the interval $[1,2]$ shows that it has an absolute maximum at $x=2$ if $p\geq 2.$
So
$(b+p)\lambda-q\geq 0\iff a\geq 4q^{2}/(2+p)^{2}.$
In light of the above inequalities, if $p$ and $q$ are positive real numbers with $p\geq 2,$ define
(5.8) $m(p,q):=\frac{4q^{2}}{(2+p)^{2}}.$
In the following several subsections, we conduct the reduction calculations on a useful partition of $[9,16).$
5.1.1. $a\in[9,10)$
We have
(5.9) $(b+2)\lambda-6\geq 0$
because $m(2,6)=9\leq a.$ In particular, $h(2)\geq 1.$ Therefore,
$h(0)>h(1)>h(2)\geq 1$
by 5.9, so
$C_{2}=(h(1)+\lambda;h(1),1^{\times 6},(\lambda-1)^{\times 3},d_{1}^{\times l_{1}},\ldots),$
with $\delta_{2}=\lambda-2,$ which implies
$C_{3}=(h(2)+\lambda;h(2),1^{\times 4},(\lambda-1)^{\times 5},d_{1}^{\times l_{1}},\ldots),$
with $\delta_{3}=\lambda-2.$
So
$C_{4}=(h(3)+\lambda||h(3),1^{\times 2},(\lambda-1)^{\times 7},d_{1}^{\times l_{1}},\ldots).$
If $h(3)\geq 1,$ then $\delta_{4}=\lambda-2,$ and
$C_{5}=(h(4)+\lambda;h(4),(\lambda-1)^{\times 9},d_{1}^{\times l_{1}},\ldots)$
because $h(3)\geq 1\implies h(4)\geq\lambda-1.$ Therefore, $\delta_{5}=2-\lambda>0.$
Now suppose $h(3)<1.$ Now $h(2)\geq 1\implies h(3)\geq\lambda-1,$ and
$1>h(3)\implies\lambda-1>h(4).$
Putting everything together, we get $1>h(3)\geq\lambda-1>h(4),$ and
$C_{4}=(h(3)+\lambda;1^{\times 2},h(3),(\lambda-1)^{\times 7},d_{1}^{\times l_{1}},\ldots),$
with $\delta_{4}=\lambda-2.$ Thus
$C_{5}=(h(4)+\lambda||(\lambda-1)^{\times 9},h(4),d_{1}^{\times l_{1}},\ldots).$
Since $\lambda-1$ exceeds $h(4)$ and $d_{1},\ldots,d_{t},$ it follows that
$\delta_{5}=h(4)+3-2\lambda=h(2)-1\geq 0,$
where the last equality follows from the last part of Lemma 5.1.
5.1.2. $a\in[10,10.24)$
In this situation, $h(2)>1,$ and by Lemma 5.2 we get
$C_{3}=(h(2)+\lambda;h(2),1^{\times 5},(\lambda-1)^{\times 5},d_{1}^{\times l_{1}},\ldots)$
with $\delta_{3}=\lambda-2,$ so
$C_{4}=(h(3)+\lambda||h(3),1^{\times 3},(\lambda-1)^{\times 7},d_{1}^{\times l_{1}},\ldots).$
If $h(3)\geq 1,$ then $\delta_{3}=\lambda-2$ so
$C_{5}=(h(4)+\lambda||h(4),1^{\times 1},(\lambda-1)^{\times 9},d_{1}^{\times l_{1}},\ldots),$
and since $h(4)\geq\lambda-1,$ we have $\delta_{5}=0.$
Otherwise, $1>h(3),$ so
$C_{4}=(h(3)+\lambda;1^{\times 3},h(3),(\lambda-1)^{\times 7},d_{1}^{\times l_{1}},\ldots),$
and $\delta_{4}=h(3)+\lambda-3=h(4)-1.$ Since $\lambda-1>h(4),$ we have $h(4)-1<0.$ Therefore,
$C_{5}=(h(3)+h(4)+\lambda-1||h(3),(\lambda-1)^{\times 7},h(4)^{\times 3},d_{1}^{\times l_{1}},\ldots).$
Since $h(3)>\lambda-1,$ we have $\delta_{5}=h(4)+1-\lambda<0.$ Therefore,
$C_{6}$ $=$ $(h(3)+2h(4)||h(3)+h(4)+1-\lambda,(\lambda-1)^{\times 5},h(4)^{\times 5},d_{1}^{\times l_{1}},\ldots)$
$=$ $(3h(3)+2\lambda-4||2h(3)-1,(\lambda-1)^{\times 5},h(4)^{\times 5},d_{1}^{\times l_{1}},\ldots).$
Now
$2h(3)-1\geq\lambda-1\iff(b+5/2)\lambda-7\geq 0\iff a\geq m(5/2,7)\approx 9.68$
Therefore, $\delta_{6}=h(3)-1<0,$ so
$C_{7}=(4h(3)+2\lambda-5||3h(3)-2,(\lambda-1)^{\times 3},h(4)^{\times 7},d_{1}^{\times l_{1}},\ldots).$
Similar to above reasoning, we have
$3h(3)-2\geq\lambda-1\iff(b+8/3)\lambda-22/3\geq 0\iff a\geq m(8/3,22/3)\approx 9.88$
Again, we have $\delta_{7}=h(3)-1,$ so
$C_{8}=(5h(3)+2\lambda-6||4h(3)-3,(\lambda-1)^{\times 1},h(4)^{\times 9},d_{1}^{\times l_{1}},\ldots).$
Now
$4h(3)-3\geq\lambda-1\iff(b+11/4)\lambda-15/2\geq 0\iff a\geq m(11/4,15/2)\approx 9.97$
[…]
$h(4)\geq.24\iff(b+4)\lambda-9.24\geq 0\iff a\geq m(4,9.24)\approx 9.49.$
Putting these estimates together, we conclude that the ordering to the right of the double bars in $C_{8}$ above is accurate and
$\delta_{8}$ $=$ $5h(3)+2\lambda-6-4h(3)+3-\lambda+1-h(4)$
$=$ $h(3)+\lambda-2-h(4)$
$=$ $h(4)-h(4)=0.$
5.1.3. $a\in[10.24,12)$
First, we note that
$h(3)=(b+3)\lambda-7>1$
for all $1\leq b\leq 2$ because $a\geq m(3,8)=10.24.$
Assume $1110.24\leq a<$ 11 . By Lemma 5.2, we have
$C_{4}=(h(3)+\lambda;h(3),1^{\times 3},(\lambda-1)^{\times 7},d_{1}^{\times l_{1}},\ldots)$
with $\delta_{4}=\lambda-2.$ Therefore,
$C_{5}=(h(4)+\lambda||h(4),1^{\times 1},(\lambda-1)^{\times 9},d_{1}^{\times l_{1}},\ldots).$
Now $h(4)>\lambda-1$ by Lemma 5.1. In particular,
$\lambda-1=\min\{1,\lambda-1,h(4)\}.$
Therefore, $\delta_{5}=0.$
Now assume $1211\leq a<$ 12 . Lemma 5.2 shows
$C_{4}=(h(3)+\lambda;h(3),1^{\times 4},(\lambda-1)^{\times 7},d_{1}^{\times l_{1}},\ldots),$
with defect $\delta_{4}=\lambda-2.$ So
$C_{5}=(h(4)+\lambda||h(4),1^{\times 2},(\lambda-1)^{\times 9},d_{1}^{\times l_{1}},\ldots).$
Assume $h(4)\geq 1.$ Recall that $h(4)\geq 1\implies h(5)\geq\lambda-1.$ Therefore,
$C_{5}=(h(4)+\lambda;h(4),1^{\times 2},(\lambda-1)^{\times 9},d_{1}^{\times l_{1}},\ldots),$
with $\delta_{5}=\lambda-2.$ Applying another Cremona transformation, we see that
$C_{6}=(h(5)+\lambda;h(5),(\lambda-1)^{\times 11},d_{1}^{\times l_{1}},\ldots),$
with defect $\delta_{6}=2-\lambda>0.$
Now suppose $h(4)<1.$ Similar to previous reasoning, we have
(5.10) $1>h(4)>\lambda-1>h(5).$
So
$C_{5}=(h(4)+\lambda;1^{\times 2},h(4),(\lambda-1)^{\times 9},d_{1}^{\times l_{1}},\ldots),$
with $\delta_{5}=\lambda-2.$ By 5.10, we get
$C_{6}=(h(5)+\lambda||(\lambda-1)^{\times 11},h(5),d_{1}^{\times l_{1}},\ldots),$
with $\delta_{6}=h(5)+3-2\lambda=h(3)-1>0,$ where the latter equality follows from the last part of Lemma 5.1.
5.1.4. $a\in[12,16)$
We start by making the following observations:
$h(5)\geq 1$ $\iff$ $a\geq m(5,12)=11.755.$
$h(6)\geq 1$ $\iff$ $a\geq m(6,14)=12.25.$
$h(7)\geq 1$ $\iff$ $a\geq m(7,16)\approx 12.642.$
If $12\leq a<13,$ then $h(5)\geq 1,$ so Lemma 5.2 implies
$C_{6}=(h(5)+\lambda;h(5),1^{\times 1},(\lambda-1)^{\times 11},d_{1}^{\times l_{1}},\ldots)$
with $\delta_{6}=0.$
If $13\leq a<14,$ then $h(6)\geq 1,$ so
$C_{7}=(h(6)+\lambda;h(6),(\lambda-1)^{\times 13},d_{1}^{\times l_{1}},\ldots),$
with $\delta_{7}=2-\lambda>0.$
For $14\leq a<15,$ we still use $h(6)\geq 1$ and Lemma 5.2 to get
$C_{7}=(h(6)+\lambda;h(6),1^{\times 1},(\lambda-1)^{\times 13},d_{1}^{\times l_{1}},\ldots),$
with $\delta_{7}=0.$
Finally, if $15\leq a<16,$ Lemma 5.1 gives
$C_{8}=(h(7)+\lambda;h(7),1^{\times 1},(\lambda-1)^{\times 14},d_{1}^{\times l_{1}},\ldots),$
with $\delta_{8}=0.$
5.2. Case II: $d_{1}>\lambda-1$
We begin this case with the following lemma.
Lemma 5.3.
We have $\lambda-1>1/2.$ In particular, if $a$ is such that $d_{1}>\lambda-1,$ and $l_{1}$ is the multiplicity of $d_{1}$ in the weight expansion of $a,$ then $l_{1}\leq 1.$
Proof.
We have $\lambda>3/2\iff 2a>9b.$ Since $a\geq 9$ and $b\leq 2,$ we have $2a\geq 18\geq 9b.$
A suitable modification of Lemma 5.2 states that under the conditions of Lemma 5.2 we have
(5.11) $C_{k+1}=(h(k)+\lambda||h(k),1^{\times\left\lfloor a\right\rfloor-(2k+1)},d_{1}^{\times l_{1}},(\lambda-1)^{\times 2k+1},\ldots)$
We handle this case on intervals of the form $[k,k+1)$ where $k$ is an integer as we did above.
5.2.1. $a\in[9,10)$
Since $h(2)\geq 1,$ Equation 5.11 gives us
$C_{3}=(h(2)+\lambda||h(2),1^{\times 4},d_{1},(\lambda-1)^{\times 5},\ldots)$
with defect $\delta_{3}=\lambda-2<0.$ So
$C_{4}=(h(3)+\lambda||h(3),1^{\times 2},d_{1},(\lambda-1)^{\times 7},\ldots).$
From this point, there are several cases to consider, each of which yields different calculations. The cases stem from how $h(3),1,$ and $d_{1}$ relate to one another. Each case below is viewed as a subcase of the assumption that $a\in[9,10).$ Such will be the case in future subsections where similar computations are demonstrated. Computations of subsequent $\delta_{k}$ will make liberal use of the identities in Lemma 5.1.
Subcase 1: $h(3)\geq 1.$ This case is the most straightforward. In this case,
$h(3)\geq 1\geq d_{1}>\lambda-1$
so $\delta_{4}=\lambda-2<0.$ Therefore,
$C_{5}=(h(4)+\lambda||h(4),d_{1},(\lambda-1)^{\times 9},\ldots).$
If $h(4)\geq d_{1},$ then $\delta_{5}=1-d_{1}\geq 0.$ If $d_{1}>h(4)>\lambda-1,$ then still $\delta_{5}=1-d_{1}\geq 0.$ Finally, if $d_{1}>\lambda-1>h(4),$ then
$\delta_{5}$ $=$ $h(4)+\lambda-d_{1}-2(\lambda-1)$
$=$ $h(4)-d_{1}+2-\lambda$
$\geq$ $h(4)-1+2-\lambda$
$=$ $h(3)-1\geq 0.$
Subcase 2: $1>h(3)>d_{1}.$ This case is very similar to the previous since, as before, we have $\delta_{4}=\lambda-2,$ and
$C_{5}=(h(4)+\lambda||h(4),d_{1},(\lambda-1)^{\times 9},\ldots).$
The analysis directly above shows that $\delta_{5}\geq 0$ in this case.
Subcase 3: $1>d_{1}>h(3).$ To begin, we need a key inequality.
Lemma 5.4.
If $d_{1}>h(3)$ and $9\leq a<10,$ we have
(5.12) $4h(3)-3d_{1}>\lambda-1>h(4)+1-d_{1}.$
Proof.
On $[9,10),$ $d_{1}=a-9,$ so the first inequality in 5.12 holds if and only if
(5.13) $(4b+11)\lambda-3a>0.$
Fix $b,$ and view the left-hand side of 5.13 as a function of $a.$ The derivative of the left-hand side is $(4b+11)\lambda^{\prime}-3,$ which is negative if and only if $\lambda^{\prime}=\frac{1}{4b}\sqrt{\frac{2b}{a}}<3/(4b+11),$ or equivalently,
$\sqrt{\frac{2b}{a}}<\frac{12b}{4b+11}.$
Now $1512b/(4b+11)\geq 12/$ 15 for $b\in[1,2],$ and $\sqrt{2b/a}\leq\sqrt{2b}/3$ on $[9,10).$ The above inequality will hold, then, if $\sqrt{2b}<36/15,$ which is certainly true. Therefore, $(4b+11)\lambda-3a$ is decreasing as a function of $a$ on $[9,10).$ Now
$\lim_{a\to 10^{-}}(4b+11)\lambda-3a=(4b+11)\sqrt{5/b}-30>0,$
where the latter inequality follows from familiar considerations using differential calculus. In summary, the function in question is positive on $[9,10),$ and the first inequality in 5.12 is hence established.
For the second inequality in 5.12, note that $\lambda-1>h(4)+1-d_{1}$ if and only if $d_{1}>h(4)+2-\lambda=h(3),$ which is given. ∎
As an immediate consequence of Lemma 5.4 and our key assumption that $d_{1}>h(3),$ we have
(5.14) $h(3)>2h(3)-d_{1}>3h(3)-2d_{1}>4h(3)-2d_{1}>\lambda-1>h(4)+1-d_{1},$
where the last inequality follows directly from the assumption that $d_{1}>h(3)$ and the familiar identity $h(3)+\lambda-2=h(4).$
These inequalities in mind, we are ready to begin reduction calculations in this case. Throughout, let $d_{2}=1-d_{1}.$ We have
$C_{4}=(h(3)+\lambda;1^{\times 2},d_{1},h(3),(\lambda-1)^{\times 7}||\ldots),$
so that $\delta_{4}=h(3)+\lambda-2-d_{1}=h(4)-d_{1}.$
If $\delta_{4}<0,$ then
$C_{5}=(h(3)+h(4)+\lambda-d_{1}||(h(4)+d_{2})^{\times 2},h(4),h(3),(\lambda-1)^{\times 7},\ldots).$
Now $1>h(3)\implies\lambda-1>h(4),$ so
$C_{5}=(h(3)+h(4)+\lambda-d_{1};h(3),(\lambda-1)^{\times 7},(h(4)+d_{2})^{\times 2},h(4)||\ldots),$
and $\delta_{5}=h(3)-d_{1}<0.$ Now $\lambda-1+h(3)-d_{1}=h(4)+d_{2}$ which, combined with 5.14, implies
$C_{6}=(2h(3)+h(4)+\lambda-2d_{1};2h(3)-d_{1},(\lambda-1)^{\times 5},(h(4)+d_{2})^{\times 4},h(4)||\ldots),$
so that again $\delta_{6}=h(3)-d_{1}<0.$ Continuing, we see
$C_{7}=(3h(3)+h(4)+\lambda-3d_{1};3h(3)-2d_{1},(\lambda-1)^{\times 3},(h(4)+d_{2})^{\times 6},h(4)||\ldots),$
so that $\delta_{7}=h(3)-d_{1}<0.$ Finally, we see that
$C_{8}=(4h(3)+h(4)+\lambda-4d_{1};4h(3)-3d_{1},(\lambda-1)^{\times 1},(h(4)+d_{2})^{\times 8},h(4),\ldots),$
and $\delta_{8}=0.$ This concludes the case where $109\leq a<$ 10 .
5.2.2. $a\in[10,11)$
Since $h(2)\geq 1,$ we start with
$C_{3}=(h(2)+\lambda;h(2),1^{\times 5},d_{1},(\lambda-1)^{\times 5}||\ldots),$
with $\delta_{3}=\lambda-2.$ So
$C_{4}=(h(3)+\lambda||h(3),1^{\times 3},d_{1},(\lambda-1)^{\times 7},\ldots).$
We claim $h(3)\geq 1$ when $d_{1}>\lambda-1.$ Indeed, $d_{1}=a-10$ on this interval, and recall [Lemma] asserts that $\lambda-1>1/2.$ So $d_{1}>\lambda-1\implies a-10>1/2,$ or equivalently, $a>10.5$ Now $h(3)\geq 1$ if and only if $(b+3)^{2}a/2b\geq 64,$ and since $(b+3)^{2}a/(2b)\geq(b+3)^{2}\cdot 10.5/(2b),$ we need only hope that the latter expression exceeds 64. If $f(x)=10.5(x+3)^{2}/2x$ on $[1,2],$ then $f$ is easily seen to be decreasing, and its least value is thus $f(2)=65.525.$ Thus $h(3)\geq 1$ as desired. As an immediate consequence, $h(4)\geq\lambda-1.$
Since $h(3)\geq 1,$ we have $\delta_{4}=\lambda-2,$ and after re-arranging the first three terms past the head, we get
$C_{5}=(h(4)+\lambda||1^{\times 1},d_{1},h(4),(\lambda-1)^{\times 9},\ldots).$
The term $h(4)$ may be greater than or equal to $d_{1}$ or less than $d_{1}$ on this interval. However, $h(4)\geq\lambda-1$ when $d_{1}>\lambda-1.$ Put another way, the least of $h(4),1,$ and $d_{1}$ must be at least as great as $\lambda-1$ on $[10,11),$ provided that $d_{1}>\lambda-1.$ In particular, $\delta_{5}=\lambda-1-d_{1}<0,$ and we get
$C_{6}=(h(5)+\lambda+1-d_{1}||h(5)+1-d_{1},\lambda-d_{1},(\lambda-1)^{\times 10},\ldots).$
If $h(5)+1-d_{1}\geq\lambda-1,$ then $\delta_{6}=d_{1}-\lambda+1>0.$ Otherwise, $h(5)+1-d_{1}<\lambda-1,$ and we have
$\delta_{6}=h(5)+3-2\lambda=h(4)+1-\lambda\geq 0.$
5.2.3. $a\in[11,12)$
Before we get started here, we claim $h(5)>d_{1}.$ First, note that $d_{1}=a-\left\lfloor a\right\rfloor=a-11.$ The desired inequality is true if and only if $(b+5)^{2}/2b\geq a,$ which holds if the former expression is at least as great as 12. The function $f(x)=(x+5)^{2}/2x$ on $[1,2]$ is decreasing, and $f(2)>12,$ so the assertion is, in fact, accurate. In particular, since $h(4)>h(5),$ it follows that $h(4)>d_{1}.$
In this case, we have
$C_{4}=(h(3)+\lambda||h(3),1^{\times 4},(\lambda-1)^{\times 7},d_{1},\ldots),$
with $\delta_{4}=\lambda-2.$ So,
$C_{5}=(h(4)+\lambda||h(4),1^{\times 2},d_{1},\ldots).$
If $h(4)\geq 1,$ we have $\delta_{5}=\lambda-2$ and
$C_{6}=(h(5)+\lambda;h(5),d_{1},(\lambda-1)^{\times 11}||\ldots)$
with $\delta_{6}=1-d_{1}\geq 0.$ If $h(4)<1,$ then
$C_{5}=(h(4)+\lambda;1^{\times 2},h(4),d_{1},(\lambda-1)^{\times 9}||\ldots),$
and $\delta_{5}=\lambda-2<0.$ Continuing, we see that
$C_{6}=(h(5)+\lambda h(5),d_{1},(\lambda-1)^{\times 11}||\ldots),$
so that $\delta_{6}=1-d_{1}\geq 0.$
5.2.4. $a\in[12,13)$
Here, we have
$C_{6}=(h(5)+\lambda||h(5),1^{\times 1},d_{1},(\lambda-1)^{\times 11},\ldots)$
so $\delta_{6}=\lambda-1-d<0.$ Then
$C_{7}=(h(6)+\lambda+1-d_{1}||h(6)+1-d_{1},\lambda-d_{1},(\lambda-1)^{\times 12},\ldots),$
with $\delta_{7}=d-\lambda+1>0.$
5.2.5. $a\in[13,14)$
Since $h(6)\geq 1,$ we get
$C_{7}=(h(6)+\lambda||h(6),d_{1},(\lambda-1)^{\times 13},\ldots),$
so that $\delta_{7}=1-d\geq 0.$
5.2.6. $a\in[14,15)$
We have $h(6)\geq 1$ so
$C_{7}=(h(6)+\lambda;h(6),1^{\times 1},d_{1},(\lambda-1)^{\times 13}||\ldots).$
Then $\delta_{7}=\lambda-1-d<0$ so that
$C_{8}=(h(7)+\lambda+1-d_{1};h(7)+1-d_{1},\lambda-d_{1},(\lambda-1)^{\times 14}||\ldots).$
In particular, $\delta_{8}=d_{1}-\lambda+1>0.$
5.2.7. $a\in[15,16)$
In this case, we have $h(7)\geq 1$ so that
$C_{8}=(h(7)+\lambda||h(7),1^{\times 1},d_{1},(\lambda-1)^{14},\ldots).$
From here, we see $\delta_{8}=\lambda-1-d_{1}<0,$ and thus
$C_{9}=(h(8)+\lambda+1-d_{1}||h(8)+1-d_{1},\lambda-d_{1},(\lambda-1)^{\times 15},...),$
and thus $\delta_{9}=d_{1}-\lambda+1>0.$